A swimming pool is 28 feet long and 21 feet wide. How many feet does a boy save by swimming diagonally across the pool instead of swimming the length and then the width of the pool?
1
Expert's answer
2012-01-19T10:32:05-0500
The diagonal length of a swimming pool is D = sqrt(28² + 21²) = sqrt(1225) = 35 feet.
So, the profit is (28+21)-35 = 49-35 = 14 feet.